OpenStudy (anonymous):

When BA=10ft, find the area of the region that is not shaded

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (jdoe0001):

I'm assume BA is the radius?

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

segment of a circle => \(\bf \cfrac{(\theta-sin(\theta)) r^2}{2}\)

OpenStudy (jdoe0001):

darn... I got that in radians.

OpenStudy (anonymous):

how do get the theta

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{(\theta-sin(\theta)) r^2}{2}\\ \cfrac{(\frac{\pi}{3}-sin\left(\frac{\pi}{3}\right)) r^2}{2} \implies \cfrac{\frac{100 \pi}{3}-100sin\left(\frac{\pi}{3}\right)}{2}\\ \cfrac{\frac{100 \pi}{3}-100\left(\frac{\sqrt{3}}{2}\right)}{2} \implies \cfrac{\frac{100 \pi}{3}-50\sqrt{3}}{2} \implies \cfrac{\frac{100\pi -150\sqrt{3}}{3}}{2}\\ \cfrac{100\pi -150\sqrt{3}}{6}\)

OpenStudy (jdoe0001):

since the equation uses radian, so I set 60 degrees to \(\bf \frac{\pi}{3}\) radians

OpenStudy (jdoe0001):

so the area that is not shaded will be the circle's full area minus the segment's \(\bf \pi r^2 -\cfrac{100\pi -150\sqrt{3}}{6}\)